You are given the following information about an economy:

Gross private domestic investment = 40

Government purchases of goods and services = 30

Gross national product (GNP) = 200

Current account balance = -20

Taxes = 60

Government transfer payments to the domestic private sector = 25

Interest payments from the government to the domesticprivate sector = 15 (Assume all interest payments by the government go to domestic households.)

Factor income received from rest of world = 7

Factor payments made to rest of world = 9

Find the following, assuming that government investment is zero:

a. Consumption

b. Net exports

c. GDP

d. Net factor payments from abroad

e. Private saving

f. Government saving

g. National saving


1
Expert's answer
2021-09-06T07:05:24-0400

GDP: GDP or gross domestic product is the estimation of the final value of all the goods and services that are produced within the domestic territory of a country during a specified period of time.


Formulation

We know that Y = C + I + G + NX  i.e GDP = Consumption expenditure + Investment + Government spending + Net export 

Thus, C = Y - I - G - NX                  ................................... eq.1

Investment I = 40

Government expenditure G = 30

 GDP Y = GNP - NFI                    ................................... eq.2

GNP = 200

Net factor income NFI  = Factor income received from rest of the world - Factor payment made to rest of the world

Net factor income NFI  = 7 - 9  = -2

Putting in the values in eq. 2 we get

GDP Y = 200 - (-2)

GDP Y = 202


calculation

Now,

Net export NX= Current account balance - Net factor income (NFI)

Net export NX= -20 - (-2)

Net export NX= -18

Putting in the values in eq 1 we get

C = 202 - 40 - 30 -(-18)

C = 150


a.

Thus the consumption in the economy is 150.


b.

The net export (NX) in the economy is -18.


c.


The GDP size of the economy is 202.


d.

The net factor payment from abroad or the NFI is -2.


e.

Private saving = (GDP - NFI - taxes + transfer payment + Interest payment) - consumption

            = (202 -(-2) - 60 +25 + 15) - 150

            = 184 - 150

Private saving = 34


f.

Government saving = (taxes - transfer payment - Interest payment) - government expenditure

                = (60 - 25 - 15) -30

                = 20 - 30

Government saving = -10


g.

National saving = Private saving + Government saving

National saving = 34 + (-10)

National saving = 24